Stan and Hilda can mow the lawn in 40 minutes if they work together. If Hilda works twice as fast as Stan, how long would it take Stan to mow the lawn alone?
I set this up as if Stan = 2x and Hilda = x
then:
40 min (1/x + 1/2x)
= 40/x + 40/2x = 40+20 = 60 x
So Hilda would take 60 minutes and Stan 120 minutes, but this doesn't seem right.
Thanks for all your help!

Answer by Paul(988)

You are correct:
Letting x be the time for individuals

1/x+1/x=1/40
2(1/x)+1/x=1/40
2/x+1/x=1/40
3/x=1/40
3=x/40
x=120

Which is the amont of time riquired for Stan to mow the lawn alone.
Paul.